# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ 5 Properties of the scaling limit under limited outcrossing
Here we provide some additional analyses of the novel scaling limits in Theorems 4.3 and 4.4, especially the coalescing random walk on the ancestral graph, motivated by their possible utility for a larger class of models than just our Moran model. For example, the simulations for the Wright-Fisher model in Figure 1 shows similar behavior as that for our Moran model in Figure 3.
@@
Observe first that the limited-outcrossing regime in Theorem 4.3 includes the case $\lambda=0$, which means that the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ (when $\lambda=0$) is equal to that of $H$ in the negligible-outcrossing regime, where $H$ is a constant random variable whose value is exponentially distributed with rate 2. Corollary 22 below covers the other extreme: as $\lambda\to\infty$, the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ converges to ${\rm Exp}(2)$, which corresponds to the partial-selfing regime with $\alpha=1$. Hence, for lineages in different individuals, the limited-outcrossing regime covers all possible cases for which $\alpha_{N}\to 1$, as we tune the magnitude of $\lambda$. This highlights the significance of the limited-outcrossing regime.
We focus our analysis on the variance of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ (for all $t\in\mathbb{R}_{>0}$) and that of the conditional expectation $\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]$. We obtain explicit formulas and their asymptotics (as $\lambda\to 0$ and $\lambda\to\infty$). The starting point of our analysis is a connection to a pair of two conditionally independent gene genealogies of the sample, in the spirit of Diamantidis et al. (2024) and Birkner et al. (2013).
@@
For any $\lambda\in\mathbb{R}_{>0}$ and $t\in\mathbb{R}_{>0}$,
@@
where $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are two conditionally independent copies of the pairwise hitting time given the same $G_{\lambda}$; they are defined in (13) with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$ and $x_{\lambda}(0)\neq y_{\lambda}(0)$.
@@
We proof only the first equality in (15) which is equivalent to $\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]=\mathbb{P}(T_{\lambda}\wedge T_{\lambda}^{\prime}>t)$. The second follows from the same argument.
@@
$=$ $\,\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t),$
where the second equality above follows from the conditional independence of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Recall from Lemma 5.1 that the random variables $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are the pairwise meeting times for two conditionally independent processes $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ with the same starting point, i.e. with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$. Each of these two processes represents the dynamics of a pair of particles on the ancestral graph $G_{\lambda}$. Note also that in Lemma 5.1 we required $x_{\lambda}\neq x_{\lambda}^{\prime}$, meaning that $T_{\lambda}$ and $T_{\lambda}^{\prime}$ apply to the case in which two individuals are sampled and two copies of the pairwise coalescence process are followed starting at these same two individuals.
In order to model the full two-pairs process, we need to consider all possible samples and all possible ancestral states of the samples. For simplicity, we combine states where possible by symmetry and restrict ourselves to a process with the following five states.
@@
Here, two dots of the same color (red or blue) correspond to two particles of the same pair; say, the two red dots correspond to $(x_{\lambda},y_{\lambda})$ and the two blue dots correspond to $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$. Two dots within a parenthesis correspond to taking the same value in $G_{\lambda}$, i.e. being in the same individual. Hence, for $i\in\{0,1,2\}$, the state $s_{i}$ corresponds to when $i$ individuals or particles in $G_{\lambda}$ contain two conditionally independent lineages.
@@
$Q_{\lambda}=\begin{pmatrix}-12&8&0&4&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2&4&0\\ 0&\lambda&-2-\lambda&0&2\\ 0&0&0&-2&2\\ 0&0&0&0&0\end{pmatrix}$ (16)
The proof of Lemma 5.2 is detailed in the Appendix. Observed that $Q_{\lambda}$ is equivalent to the corresponding matrices for the Kingman coalescent with recombination between two loci (Kaplan and Hudson, 1985; Pluzhnikov and Donnelly, 1996; Simonsen and Churchill, 1997); see also Wakeley (2009, Chapter 7) and Durrett (2008, Chapter 3). More precisely, our model here becomes that in Wakeley (2009, Figure 7.7) if we divide all our rates by two, and then make the substitution $\lambda/2\to\rho$. This $1/2$ on $\lambda$ comes from the fact that each splitting event is only observed with probability $1/2$.
@@
$\mathrm{Cov}_{s_{0}}{\left(T_{1},T_{2}\right)}$ $=\frac{4}{\lambda^{2}+26\lambda+72}$ (17)
$\mathrm{Cov}_{s_{1}}{\left(T_{1},T_{2}\right)}$ $=\frac{6}{\lambda^{2}+26\lambda+72}$ (18)
$\mathrm{Cov}_{s_{2}}{\left(T_{1},T_{2}\right)}$ $=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (19)
@@
The first equality follows directly from (15). Since $(T_{\lambda},T_{\lambda}^{\prime})=(T_{1},T_{2})$ given state $s_{2}$, (19) is precisely the second equality in (20). ∎
In the Discussion, we will consider the significance of (20) for an understanding how coalescence times depend on the pedigree, represented in this case by $G_{\lambda}$.
Next, we consider the variance of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$. We shall compute the second moment of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ in Lemma 5.7 and use it to show the following.
@@
In particular, for any $t\in\mathbb{R}_{>0}$,
$\mathbb{P}(T_{\lambda}>t|G_{\lambda})\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}\textbf{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{ as }\lambda\to 0\\ e^{-2t}&\quad\text{ as }\lambda\to\infty\end{dcases}$ (22)
Lemma 5.7 and Proposition 5.4 will be proved in Section 5.1.
We observe that the covariance $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is connected to the error between the conditional survival function $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and the unconditional survival function $\mathbb{P}(T_{\lambda}>t)=e^{-2t}$.
@@
By Lemma 5.1, the LHS is equal to
@@
In agreement with previous work—see equations (3) and (52) in Simonsen and Churchill (1997)—from (19) and the fact $\mathrm{Var}{\left(T_{\lambda}\right)}=1/4$, we see that the correlation coefficient $\mathrm{Corr}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is exactly the same as $\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$. Equivalently, $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}=\frac{1}{4}\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$, and it is also equal to the left hand side of (23).
Depending on $\lambda\in(0,\infty)$, limited outcrossing displays a range of behaviors between the two qualitatively different extremes of negligible outcrossing and partial selfing. Figure 8 illustrates this using the law of total variance for the coalescence time of two gene copies sampled from two different individuals. Let $T$ and $G$ represent this “diff” coalescence time and its associated ancestral graph, for simplicity dropping the subscript $\lambda$ we attached to these previously in Section 5. Then we can write
$\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ (24)
and note that $\textrm{Var}{\left(T\right)}=1/4$, because $T$ by itself is exponentially distributed with rate parameter $2/(2-1)=2$. The second term on the right-hand side of (24), $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, is given by (20). The first term on the right-hand side of (24) is just $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}=1/4-\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, but both terms are displayed in Figure 8 for the sake of illustration.
@@
When $\lambda$ is very small, on the far left in Figure 8, variation in the mean coalescence time among graphs, $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, accounts for the bulk of variation in $T$. Correspondingly, there is little variation in coalescence times given the graph. Figure 8 displays the expectation of the latter, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$, but $\textrm{Var}{\left(T|G\right)}$ itself will be small on any graph in this case. With probability approaching one as $\lambda\to 0$, the individuals containing the two ancestral lineages will experience only selfing events in their ancestries, up to and including the time they descend from a common ancestral individual. Pairwise coalescence times at every locus in the genome starting from these two individuals will converge on the time of this ancestor and $\textrm{Var}{\left(T|G\right)}$ should be minimal. Variation in the waiting time to this first common ancestral individual among graphs will be the primary source of variation in $T$. In the limit $\lambda\to 0$, this waiting time is exponentially distributed with rate parameter $2$, so $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ approaches $\textrm{Var}{\left(T\right)}=1/4$. It is given by a Kingman coalescent process made effectively haploid by selfing, just as under negligible outcrossing.
@@
When $\lambda$ is very large, on the far right in Figure 8, single ancestral graphs do little to constrain $T|G$. These graphs are large. In their construction backward in time, the number of ancestral lineages grows quickly to reach a quasi-stable distribution, specifically a zero-truncated Poisson distribution with mean $\lambda$, as Mano (2009) showed for the ancestral recombination graph. The waiting time to the event that completes the graph, when there is just one lineage left, will be long, $\sim 2e^{\lambda}/\lambda^{2}$ in expectation using equation (1.1) in Griffiths and Marjoram (1997). Given the graph, each of the pair of genetic lineages starting from the two sampled individuals will trace back through a large number of splitting events before the two meet in a single individual and coalesce. As $\lambda\to\infty$, all variation in $T$ will be due to this random process of coalescence given the graph. Corollary 22 shows that this conditional coalescent process converges on that of the partial-selfing model with $\alpha_{N}\to 1$. Proposition 5.5 shows further that $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ in (20) is a key measure (in $L^{2}$) of the distance of the CDF under limited outcrossing to that under the corresponding model of partial selfing.
@@
To summarize what is shown in Figure 8, limited outcrossing becomes like negligible outcrossing as $\lambda\to 0$ and like partial selfing as $\lambda\to\infty$. With respect to the pre-limiting Moran model, the results and intuitions about Figure 8 concern the behavior at the boundary of complete or nearly complete selfing ($\alpha_{N}\to 1$). All at this boundary, we have: first negligible outcrossing, then limited outcrossing, then an extreme case of partial selfing with $\alpha=1$. The entire rest of the range of selfing rates $\alpha\in[0,1]$ is covered by the partial-selfing model.
5.1 Proof for Proposition 5.4
Observe that the two pairs of particles $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ absorbed at $T_{\lambda}\wedge T_{\lambda}^{\prime}$ is a continuous-time Markov chain, with state space $\{s_{0},s_{1},s_{2},s_{\Delta}\}$, where we collapsed $s_{\Delta,2}$ and $s_{\Delta,2}$ into a single state called $s_{\Delta}$. This absorbed process, called $\widetilde{S}=(\widetilde{S}_{t})_{t\in\mathbb{R}_{+}}$, has 4 states and has transition rate matrix $R_{\lambda}$ obtained from (16) by collapsing $s_{\Delta,2}$ and $s_{\Delta,2}$ into a single state $s_{\Delta}$. By Lemma 5.2, the process $\widetilde{S}$ has rate matrix
$R_{\lambda}=\begin{pmatrix}A_{\lambda}&{\bf v}^{\,\rm T}\\ {\bf 0}&0\end{pmatrix},$ (25)
where $A_{\lambda}$ is the matrix (27) and ${\bf v}=\begin{pmatrix}4&4&2\end{pmatrix}$. Note that
$T_{\lambda}\wedge T_{\lambda}^{\prime}=\inf\{t\in\mathbb{R}_{+}:\,S_{t}\in\{s_{\Delta,1},s_{\Delta,2}\}\}=\inf\{t\in\mathbb{R}_{+}:\,\widetilde{S}_{t}=s_{\Delta}\}$ (26)
is the first time the process $\widetilde{S}$ reaches $s_{\Delta}$. The joint distribution of $T_{\lambda}$ and $T_{\lambda}^{\prime}$ can be computed as in (Diamantidis et al., 2024, Lemma A1) but we omit this here since it is not needed.
We now compute the second moment (and hence also the variance) of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
Lemma 5.7.
For any $\lambda\in\mathbb{R}_{>0}$ and $t\in\mathbb{R}_{>0}$, $\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]$ is equal to the sum of the entries of the last row of the 3 by 3 matrix $e^{tA_{\lambda}}$, where
$A_{\lambda}=\begin{pmatrix}-12&8&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2\\ 0&\lambda&-2-\lambda\end{pmatrix}.$ (27)
Proof of Lemma 5.7. We first show that
$\mathbb{E}[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}]=\begin{pmatrix}0&0&1&0\end{pmatrix}e^{tR_{\lambda}}\begin{pmatrix}1&1&1&0\end{pmatrix}^{T}.$ (28)
Recall the absorbed chain $\widetilde{S}_{\lambda}=(\widetilde{S}_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ mentioned earlier in this section. Since $x_{\lambda}(0)=x_{\lambda}^{\prime}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)$, we have that the initial condition $\widetilde{S}_{\lambda}(0)=s_{2}$. Then for any $t\in\mathbb{R}_{+}$, by (15),
$\mathbb{E}\left[\mathbb{P}(T_{\lambda}>t|G_{\lambda})^{2}\right]=$ $\,\mathbb{P}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t)$
$=$ $\,\mathbb{P}_{s_{2}}(\widetilde{S}_{\lambda}(t)\in\{s_{0},s_{1},s_{2}\})$
$=$ $\,(0,0,1,0)\,e^{tR_{\lambda}}\,(1,1,1,0)^{T},$
where the second last equality follows from (26). Hence (28) is established. The right hand side of (28) is the sum of the entries of the last row of the 3 by 3 matrix $e^{tA_{\lambda}}$. Hence Lemma 5.7 follows. ∎
Proof of Proposition 5.4. We will employ equation (28) in the proof of Lemma 5.7. We begin by establishing the rate of convergence as $\lambda\to 0$. We write $R_{\lambda}=C+D\lambda$, where
$C:=\begin{pmatrix}-12&8&0&4\\ 0&-6&2&4\\ 0&0&-2&2\\ 0&0&0&0\end{pmatrix}\quad\text{and}\quad D:=\begin{pmatrix}0&0&0&0\\ \frac{1}{2}&-\frac{1}{2}&0&0\\ 0&1&-1&0\\ 0&0&0&0\end{pmatrix}.$ (29)
Then, using for instance Higham (2008, eqn. 10.40),

### v2

@@ 5 Additional properties of limited outcrossing
Here we provide further analyses and illustrations of the novel scaling limits in Theorems 4.1 and 4.5, in particular the coalescing random walk on the ancestral graph. We note that the simulations for the Wright-Fisher model in Figure 1 show similar behavior as those for our Moran model in Figure 3.
@@
Observe first that the limited-outcrossing regime in Theorem 4.1 includes the case $\lambda=0$, which means that the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ when $\lambda=0$ is equal to that expected under negligible outcrossing, namely a single draw from an exponential distribution with rate 2. The other extreme, $\lambda\to\infty$, is covered by (19) below. Namely, as $\lambda\to\infty$, the distribution of $T_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda})$ converges to ${\rm Exp}(2)$, which corresponds to the partial-selfing regime with $\alpha=1$. Hence, for lineages in different individuals, the limited-outcrossing regime covers all possible cases for which $\alpha_{N}\to 1$, as we tune the magnitude of $\lambda$. This highlights the significance of the limited-outcrossing regime.
To understand how information relevant to conditional gene genealogies can persist even in the limit $N\to\infty$, we introduce a discrete-time ancestral graph, $G^{N}=\left(G^{N}(k)\right)_{k\in\mathbb{Z}_{+}}$, which is a subset of the pedigree containing all potential individuals through which our sample of $n=2$ gene copies may have been transmitted. Each realization of $G^{N}$ is constructed by the following two rules. We refer to potential ancestral individuals of the sample as nodes. When a node is the result of an outcrossing event, it fragments into two nodes. When one node is the offspring of another node in a selfing event, the two coagulate into a single node. Individual genetic lineages are transmitted through this graph according to Mendel’s laws as in Section 2. A realization of $G^{N}$ is shown in Figure 5.
Figure 5: A realization of $G^{N}$ containing potential ancestors of the two sampled individuals (ovals) at the bottom. In past time-step $3$, the ancestral individual on the left undergoes an outcrossing event. In past time-step $4$, one of the nodes from this event is the offspring of one of the potential ancestral individuals of the sampled individual on the right, and these two nodes coagulate. In the fifth time-step, the right most particle experiences an outcrossing in which exactly one of the parents is a node in the graph. This may appear in the discrete-time ancestral graph but will not occur in the limiting ancestral graph.
This graphical structure will be poorly behaved as $N\to\infty$ if the number of nodes diverges as the population size goes to infinity. For the random variable $G^{N}$ to have a suitable limit as the population size goes to infinity we require that the rate at which nodes fragment, approximately $(1-\alpha_{N})N^{-1}$, and the rate at which pairs of nodes coagulate, of order $N^{-2}$, are comparable. This comparability is taken to mean that the ratio of these two rates, $N(1-\alpha_{N})$, converges as $N$ goes to infinity to $\lambda\in\mathbb{R}_{+}$. This reasoning gives rise to the following lemma.
@@
Suppose $N(1-\alpha_{N})\to\lambda\in\mathbb{R}_{+}$. The discrete-time ancestral graph under the coalescence timescale $\left(G^{N}(\lfloor tN^{2}\rfloor)\right)_{t\in\mathbb{R}_{+}}$ converges in distribution to the ancestral graph $G_{\lambda}$ as $N\to\infty$.
The proof of Lemma 5.1 is given in the Appendix, Section 7.2.2. This requires additional details which we have omitted here for simplicity. The significance of Lemma 5.1 is that the simple subgraph $G^{N}$ captures all the relevant information in the pedigree for the limiting ancestral graph $G_{\lambda}$. One difference between these graphs is that ancestral genetic lineages do not coalesce immediately at each overlap event in $G^{N}$, whereas they do necessarily at $T_{\lambda}$ in the limiting graph $G_{\lambda}$.
Next, we consider the variances of the conditional survival probability $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and of the conditional expectation $\mathbb{E}\left[T_{\lambda}|G_{\lambda}\right]$. We obtain explicit formulas and their asymptotics as $\lambda\to 0$ and $\lambda\to\infty$. The starting point of our analysis is a pair of conditionally independent gene genealogies given the graph for the sampled individuals, cf.  Diamantidis et al. (2024) and Birkner et al. (2013). Figure 6 depicts realizations of coalescence times for two such loci.
Lemma 5.2.
For any $\lambda\in(0,\infty)$ and $t\in(0,\infty)$,
@@
where $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are two conditionally independent copies of the pairwise hitting time given the same $G_{\lambda}$, as defined in (10) with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$ and $x_{\lambda}(0)\neq y_{\lambda}(0)$.
@@
First, note that $\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t|G_{\lambda})$ has mean $\mathbb{P}_{x_{\lambda}(0)\neq y_{\lambda}(0)}(T_{\lambda}>t)=e^{-2t}$ which is the same for all $\lambda$. The last equality holds because every pair of particles in the ancestral graph $G_{\lambda}$ coagulates at rate $2$. Therefore, the first equality in (12) follows from the following equality:
@@
$=$ $\,\mathbb{P}_{x_{\lambda}^{\prime}(0)=x_{\lambda}(0)\neq y_{\lambda}(0)=y_{\lambda}^{\prime}(0)}(T_{\lambda}\wedge T^{\prime}_{\lambda}>t).$
The second equality in (12) follows from the conditional independence of $T_{\lambda}$ and $T_{\lambda}^{\prime}$. ∎
Figure 6: Two conditionally independent realizations of the pairwise process, $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$. The hitting times of these random walks are marked by a purple dot. These times correspond to $T_{\lambda}$ and $T_{\lambda}^{\prime}$, the coalescence times of two unlinked loci given the same pedigree. In each realization, $x_{\lambda}$ is in blue and $y_{\lambda}$ in red. At each splitting event in the graph, where a particle splits into two, each walk chooses which particle path to follow fairly, i.e. with probability $1/2$ each. With respect to $\{t_{1},t_{2},t_{3}\}$ shown in Figure 4, here $T_{\lambda}=t_{1}$ on the left and $T_{\lambda}^{\prime}=t_{2}$ on the right.
The random variables $T_{\lambda}$ and $T_{\lambda}^{\prime}$ in Lemma 5.2 are the pairwise meeting times for two conditionally independent processes $(x_{\lambda},y_{\lambda})$ and $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$ with the same starting point, i.e. with $(x_{\lambda}(0),\,y_{\lambda}(0))=(x_{\lambda}^{\prime}(0),\,y_{\lambda}^{\prime}(0))$. They correspond to the ancestral processes for $n=2$ at two unlinked loci starting from the same two sampled individuals, as in Figure 6. To model the full two-pairs process, we consider all possible possible ancestral states of the samples. For simplicity, we combine states where possible by symmetry and restrict ourselves to a process with the following five states.
@@
Here, two dots of the same color (blue or red) correspond to two particles of the same pair; say, the two blue dots correspond to $(x_{\lambda},y_{\lambda})$ and the two red dots correspond to $(x_{\lambda}^{\prime},y_{\lambda}^{\prime})$. Two dots in parentheses correspond to taking the same value in $G_{\lambda}$, i.e. being in the same individual. Hence, for $i\in\{0,1,2\}$, the state $s_{i}$ corresponds to when $i$ individuals or particles in $G_{\lambda}$ contain two conditionally independent lineages.
@@
$Q_{\lambda}=\begin{pmatrix}-12&8&0&4&0\\ \frac{\lambda}{2}&-6-\frac{\lambda}{2}&2&4&0\\ 0&\lambda&-2-\lambda&0&2\\ 0&0&0&-2&2\\ 0&0&0&0&0\end{pmatrix}.$ (13)
[…]
@@
$\mathrm{Cov}_{s_{0}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{4}{\lambda^{2}+26\lambda+72},$ (14)
$\mathrm{Cov}_{s_{1}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{6}{\lambda^{2}+26\lambda+72},$ (15)
$\mathrm{Cov}_{s_{2}}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ $=\frac{1}{2}\frac{\lambda+36}{\lambda^{2}+26\lambda+72}.$ (16)
@@
The first equality in (17) follows directly from (12), the second from (16). ∎
Next, we consider asymptotics of the variance of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$.
@@
This follows from the computation of the second moment of $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ in Lemma 7.23 in the Appendix. Then we have, for any $t\in(0,\infty)$,
$\mathbb{P}(T_{\lambda}>t|G_{\lambda})\to\begin{dcases}\textbf{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{ as }\lambda\to 0\\ e^{-2t}&\quad\text{ as }\lambda\to\infty\end{dcases}.$ (19)
The proofs of both Lemma 7.23 and Proposition 5.5 are given in the Appendix, Section 7.3.
We observe that the covariance $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is connected to the difference between the conditional survival function $\mathbb{P}(T_{\lambda}>t|G_{\lambda})$ and the unconditional survival function $\mathbb{P}(T_{\lambda}>t)=e^{-2t}$.
@@
By Lemma 5.2, the left-hand side is equal to
@@
In agreement with previous work—see (3) and (52) in Simonsen and Churchill (1997)—from (16) and the fact $\mathrm{Var}{\left(T_{\lambda}\right)}=1/4$, we see that $\mathrm{Corr}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}$ is exactly the same as $\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$. Equivalently, $\mathrm{Cov}{\left(T_{\lambda},T_{\lambda}^{\prime}\right)}=\frac{1}{4}\mathbb{P}(T_{\lambda}=T_{\lambda}^{\prime})$, and it is also equal to the left hand side of (20). The covariance formula obtained here is consistent with Wilton et al. (2015, Eq. 9) by taking $2T_{\lambda}=T_{1}$ and $2T_{\lambda}^{\prime}=T_{2}$; the rescaling is due to the fact that the marginals on $T_{\lambda}$ and $T_{\lambda}^{\prime}$ are exponentially distributed with rate $2$, while those of $T_{1}$ and $T_{2}$ are with rate $1$.
Depending on $\lambda\in(0,\infty)$, limited outcrossing displays a range of behaviors between the two qualitatively different extremes of negligible outcrossing and partial selfing. Figure 7 illustrates this using the law of total variance for the coalescence time of two gene copies sampled from two different individuals. Let $T$ and $G$ represent this “diff” coalescence time and its associated ancestral graph, for simplicity dropping the subscript $\lambda$ we attached to these previously. Then we can write
$\textrm{Var}{\left(T\right)}=\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}+\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}.$ (21)
Note that $\textrm{Var}{\left(T\right)}=1/4$, because $T$ by itself is exponentially distributed with rate parameter $2/(2-1)=2$. The second term on the right-hand side of (21), $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, is given by (17). The first term on the right-hand side of (21) is just $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}=1/4-\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, but both terms are displayed in Figure 7 for the sake of illustration.
@@
When $\lambda$ is very small, on the far left in Figure 7, variation in the mean coalescence time among graphs, $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$, accounts for the bulk of the total variation in $T$. Correspondingly, there is little variation in coalescence times given the graph. Figure 7 displays the expectation of the latter, $\mathbb{E}{\left[\textrm{Var}{\left(T|G\right)}\right]}$, but $\textrm{Var}{\left(T|G\right)}$ itself will be small on any graph in this case. With probability approaching one as $\lambda\to 0$, the individuals containing the two ancestral lineages will experience only selfing events in their ancestries, up to and including the time they descend from a common ancestral individual. Pairwise coalescence times at every locus in the genome starting from these two individuals will converge on the time of this ancestor and $\textrm{Var}{\left(T|G\right)}$ should be minimal. Variation in the waiting time to this first common ancestral individual among graphs will be the primary source of variation in $T$. In the limit $\lambda\to 0$, this waiting time is exponentially distributed with rate parameter $2$, so $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ approaches $\textrm{Var}{\left(T\right)}=1/4$. It is given by a Kingman coalescent process made effectively haploid by selfing, just as under negligible outcrossing.
@@
When $\lambda$ is very large, on the far right in Figure 7, single ancestral graphs do little to constrain $T|G$. These graphs are large. In their construction backward in time, the number of ancestral lineages grows quickly to reach a quasi-stable distribution, specifically a zero-truncated Poisson distribution with mean $\lambda$, as Mano (2009) showed for the ancestral recombination graph. The waiting time to the event that completes the graph, when there is just one lineage left, will be long, $\sim 2e^{\lambda}/\lambda^{2}$ in expectation using equation (1.1) in Griffiths and Marjoram (1997). Given the graph, each of the pair of genetic lineages starting from the two sampled individuals will trace back through a large number of splitting events before the two meet in a single individual and coalesce. As $\lambda\to\infty$, all variation in $T$ will be due to this random process of coalescence given the graph. Equation (19) shows that this conditional coalescent process converges on that of the partial-selfing model with $\alpha_{N}\to 1$. Proposition 5.6 shows further that $\textrm{Var}{\left(\mathbb{E}{\left[T|G\right]}\right)}$ in (17) is a key measure of the distance of the CDF under limited outcrossing to that under the corresponding model of partial selfing.
@@
To summarize what is shown in Figure 7, limited outcrossing becomes like negligible outcrossing as $\lambda\to 0$ and like partial selfing as $\lambda\to\infty$. With respect to the pre-limiting Moran model, the results and intuitions about Figure 7 concern the behavior at the boundary of complete or nearly complete selfing ($\alpha_{N}\to 1$). All at this boundary, we find: first negligible outcrossing, then limited outcrossing, then an extreme case of partial selfing with $\alpha=1$. The entire rest of the range of selfing probabilities $\alpha\in[0,1]$ is covered by the partial-selfing model.
Finally, although a treatment of recombination is beyond the scope of this work, it is relatively straightforward to describe the joint ancestral process of a sample of size two at two linked loci under limited outcrossing. That is, we can extend the analysis giving the rate matrix (13) and the covariances (14)-(16) to the case of two linked loci.